how do i find the square root of 6.25
Found 3 solutions by Alan3354, solver91311, richwmiller:
Answer by Alan3354(69443)

Using a calculator is one way.
= 2.5, and -2.5

Answer by solver91311(24713)



Take out your calculator. Press the "6", press the ".", press the "2", press the "5", press the square root key, read the display.

Or, recall that 25 times 25 is 625, so 2.5 times 2.5 must be 6.25.


John


Answer by richwmiller(17219)

20^2=400
30^2=900
sqrt(625) =+-25
2^2=4
3^2=9
sqrt(6.25)=+-2.5